OpenStudy (anonymous):

Which of the following is the slope between the two points (5, 1)and (4, 6)?

myininaya (myininaya):

\[m=\frac{y_2-y_1}{x_2-x_1}=\frac{6-1}{4-5}\]

OpenStudy (anonymous):

what?

myininaya (myininaya):

what's wrong?

OpenStudy (anonymous):

the test gives me 1 over 5 an -1 over 5 then 5 and -5 for answers

myininaya (myininaya):

ok you can simplify what I have

OpenStudy (anonymous):

i dont know how to -.-

myininaya (myininaya):

6-1=what?

OpenStudy (anonymous):

5

myininaya (myininaya):

right! now what is 4-5=?

OpenStudy (anonymous):

-1

myininaya (myininaya):

so what is 5/(-1)=?

OpenStudy (anonymous):

1/-5

myininaya (myininaya):

no the 5 is on top not on bottom

OpenStudy (anonymous):

it doesnt have that as an answer

myininaya (myininaya):

5/(-1)=?

myininaya (myininaya):

do you know 5(-1)=?

OpenStudy (anonymous):

oh - 5

myininaya (myininaya):

right! 5/(-1) is the same as saying 5(-1) which is -5

OpenStudy (anonymous):

so -5= answer!!!